OpenStudy (anonymous):

Needs Checking 5x/2x * (x-4)/(x-1) 5x(x-4)=5x^2-20x 2x(x-1)=2x^2-2x -20/-2=10/1 x^2 cancel out leaving me with 10x+5 over 2

OpenStudy (loser66):

\[\dfrac{5x}{2x}*\dfrac{x-4}{x-1}\] Right?

OpenStudy (anonymous):

Correct

OpenStudy (loser66):

\[\dfrac{5\cancel{x}}{2\cancel{x}}*\dfrac{x-4}{x-1}\\=\dfrac{5(x-4)}{2(x-1)}\]

OpenStudy (anonymous):

Loser is correct

OpenStudy (anonymous):

Lol loser.

OpenStudy (anonymous):

So I took the long way.. and Ouch Opi ;)

OpenStudy (loser66):

I am a BIG Loser as @LucasPsch said, not a small loser as @OpiGeode stated, :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

You seem hot. Got any social networking?